OpenStudy (anonymous):

Solve the following system of equations by using the substitution method. x – y = 3 2x + 2y = 2 Answer (3, 4) (-4, 5) (2, -1) Inconsistent

OpenStudy (anonymous):

the answer is (2,-1)

OpenStudy (anonymous):

we have a system !: So look : -2x+2y=-6 2x+2y=2 (we had *-2 caz want to kill x s with sum the equations !) so now sum : have : 4y = -4 y = -1 and x : x-y=3 X+1=3 So x is 2 and have : (2 , -1 )

OpenStudy (anonymous):

Got it ?!:)

OpenStudy (anonymous):

if x=3-y..........(1) by using substitution let 2x+2y=2.........(2) putting (1) in (2) you have 2(3+y)+2y=2 where y=-1 then substituting y back in (1) you have x=2

OpenStudy (anonymous):

@nelotsak @E.ali Thank you both I was trying to solve a different problem while waiting for someone to help . . . .Thank you both again for the help (: and your well explanations